# Two dimensional dimers beyond planarity

### v2

@@ 4 Reflection positivity and its consequences
In this section we show that the complex measure $\langle\,\,\,\rangle_{L,K,\beta,h,r}$ which has been introduced in Section 2 is reflection positive and use this property to derive convexity and monotonicity properties for the two-point function, following [23], and to derive a Chessboard estimate, which is a well known consequence. The reflection positivity property of the complex measure reflects an equivalent reflection positivity property of the double dimer model [30]. It relies on the complex duality relation between vertices having opposite parity in (2.4) and its proof is classical [32, Chapter 10].
@@ 4.1 Reflection positivity
We let $\mathcal{S}$ be a plane which is orthogonal to the Cartesian vector $e_{i}$ for some $i\in\{1,2,3\}$ and which intersects the midpoint of some edges. We let $\Theta:\mathbb{S}_{L,K}\mapsto\mathbb{S}_{L,K}$ be a reflection operator with respect to such plane.
Let us identify two (disjoint, connected) torus halves, $\mathbb{S}_{L,K}^{\pm}$, such that $\Theta(\mathbb{S}_{L,K}^{\pm})=\mathbb{S}_{L,K}^{\mp}$. By a slight abuse of notation we keep using $\Theta:\Omega_{s}\mapsto\Omega_{s}$ to denote a reflection operator that associates to each configuration $\boldsymbol{s}=(s_{x})_{x\in V}$ the reflected spin configuration $\Theta(\boldsymbol{s})=(s_{\Theta(x)})_{x\in\mathbb{S}_{L,K}}$.
@@
$\forall\boldsymbol{s}\in\Omega_{s}\quad\Theta f(\boldsymbol{s}):={f(\Theta(\boldsymbol{s}))}.$
We say that the function $f:\Omega_{s}\mapsto\mathbb{C}$ has domain $A\subset\mathbb{S}_{L,K}$ if for each pair of configurations $\boldsymbol{s},\boldsymbol{s}^{\prime}$ such that $s_{x}=s_{x}^{\prime}$ for each $x\in A$, $f(\boldsymbol{s})=f(\boldsymbol{s}^{\prime})$. Let $\mathcal{A}^{\pm}$ be the set of functions $f:\Omega_{s}\mapsto\mathbb{C}$ which have domain $\mathbb{S}_{L,K}^{\pm}$,.
Proposition 4.1.
@@
This proof strategy is classical and can be found in several papers about reflection positivity, see for example [32, Chapter 10]. We observe that the property (i) simply follows by symmetry, hence we turn to the proof of (ii). Let us define for any measurable function $f:\Omega_{s}\mapsto\mathbb{C}$ the operator
[…]
and define $\Omega_{\pm}=S^{\mathbb{S}^{\pm}_{L,K}}$. By symmetry we have that
[…]
Now let $E^{\pm}$ be the subset of of edges of $\mathbb{S}_{L,K}$ which are entirely contained in $\mathbb{S}_{L,K}^{\pm}$ and $E_{R}$ be the set of edges intersecting both $\mathbb{S}_{L,K}^{-}$ and $\mathbb{S}_{L,K}^{+}$. We observe that we can write
$\langle f\,\,\overline{\Theta f}\rangle_{L,K,\beta,h,r}=\langle f\,\,\overline{\Theta f}\,\,e^{\mathcal{H}_{+}}e^{\mathcal{H}_{-}}e^{\sum_{\{x,y\}\in E_{R}}\beta e^{is_{x}^{1}}\,\,\Theta(e^{is_{x}^{1}})}e^{\sum_{\{x,y\}\in E_{R}}\beta e^{-is_{x}^{2}}\,\,\Theta(e^{-is_{x}^{2}})}\rangle_{0}$ (4.3)
@@
±$\mathcal{H}_{\pm}(\boldsymbol{s}):=\sum\limits_{\{i,j\}\in E^{\pm}}e^{i(s_{i}^{1}-s_{j}^{1})}+\sum\limits_{\{i,j\}\in E^{\pm}}e^{-i(s_{i}^{2}-s_{j}^{2})}+\,h\sum\limits_{k\in\mathbb{S}_{L,K}^{\pm}}\,2\,\cos(rs^{1}_{k})$
and we used the convention that if $\{x,y\}\in E_{R}$, then $x\in\mathbb{S}_{L,K}^{+}$ and $y\in\mathbb{S}_{L,K}^{-}$. Noting that $e^{\mathcal{H}_{\pm}}=\overline{\Theta e^{\mathcal{H}_{\mp}}}$ and expanding the last two exponentials in (4.3) we obtain that
$\langle f\,\,\overline{\Theta f}\rangle_{L,K,\beta,h,r}\\ =\sum\limits_{m^{1},m^{2}\in{\mathbb{N}_{0}}^{E_{R}}}\Big{\langle}f\,\,\overline{\Theta f}\,\,e^{\mathcal{H}_{+}}\,\Theta\overline{\big{(}e^{\mathcal{H}_{+}}\big{)}}\prod_{\{x,y\}\in E_{R}}\frac{\big{(}\beta e^{is_{x}^{1}}\,\,\Theta(\overline{e^{is_{x}^{1}}})\Big{)}^{m_{\{x,y\}}^{1}}}{m_{\{x,y\}}^{1}!}\frac{\big{(}\beta e^{-is_{x}^{2}}\,\,\Theta(\overline{e^{-is_{x}^{2}}})\Big{)}^{m_{\{x,y\}}^{2}}}{m^{2}_{\{x,y\}}!}\Big{\rangle}_{0}.$ (4.4)
For each term in the previous sum we set $h=fe^{\mathcal{H}_{+}}\prod_{\{x,y\}\in E_{R}}e^{im^{1}_{\{x,y\}}s_{x}^{1}}e^{-im^{2}_{\{x,y\}}s_{x}^{2}}.$ We then note that $h\in\mathcal{A}^{+}$ and that $\langle f\,\,\overline{\Theta f}\rangle_{L,K,\beta,h,r}$ can be expressed as a sum of terms of the form $\langle h\,\,\overline{\Theta(h)}\rangle_{0}$ multiplied by positive constants. By applying (4.2) for each term in the previous sum we then deduce (ii), concluding the proof. ∎
@@ 4.2 Convexity and monotonicity of two-point functions
The vector of functions $F=(F_{x})_{x\in\mathbb{S}_{L,K}}$, with $F_{x}:S\mapsto\mathbb{R}$, is said to be reflection invariant if for any $x,y,z\in\mathbb{S}_{L,K}$ the two following conditions hold
@@
    $\Theta F_{x}=F_{\Theta x}$ for any reflection through edges $\Theta$,
@@
    $\langle F_{x}F_{y}\rangle=\langle F_{x+z}F_{y+z}\rangle\geq 0$, where the sum is with respect to the torus metric.
The next lemma will allow us to derive convexity properties for two-point functions, such us the functions in the RHS of (3.14) and (3.15).
@@
Consider any reflection plane $\mathcal{S}$, let $\Theta$ be the corresponding reflection operator and $\mathbb{S}_{L,K}^{\pm}\subset\mathbb{S}_{L,K}$ the corresponding torus halves as in Section 4.1. Let $(F_{x})_{x\in\mathbb{S}_{L,K}}$ be a reflection invariant vector of functions such that $F_{o}$ has domain $D\subset\mathbb{S}_{L,K}$, with $o\in D$, and define $\ell:=\sup\{d(o,x):x\in D\}$. Let $Q\subset\mathbb{S}_{L,K}$, $Q^{\pm}:=(Q\cap\mathbb{S}_{L,K}^{\pm})\cup\Theta(Q\cap\mathbb{S}_{L,K}^{\pm})$, assume that $|Q^{+}|=|Q^{-}|$ and that $d_{H}(Q^{\pm},\mathcal{S})>\ell$, where $d_{H}$ is the Hausdorff distance with respect to the L2 norm. Then,
@@
$\mathcal{G}(x_{n})\geq\mathcal{G}(x_{n+2})\quad\quad\mathcal{G}(x_{k})\geq\mathcal{G}(x_{k+2}).$
For the proof of the proposition (given the analogous of Proposition 4.1) we refer to [23] or to [24, Theorem 2.1].
@@ 4.3 Chessboard estimates
$f^{t}:=\begin{cases}\overline{\Theta_{k}\circ\Theta_{k-1}\,\ldots\,\circ\Theta_{1}\,(f)}&\mbox{ if $k$ is even,}\\ \Theta_{k}\circ\Theta_{k-1}\,\ldots\,\circ\Theta_{1}\,(f)&\mbox{ if $k$ is odd.}\end{cases}$
@@
Let $f=(f_{t})_{t\in\mathbb{S}_{L,K}}$ be a sequence of good functions having domain $\{o\}$. We have that,
@@
Given Proposition 4.1, the proof of Propoposition 4.5 is classical and is based on performing a sequence of reflections, we refer for example to [32, Theorem 10.11] for its proof.
@@ 5 Probabilistic estimates
In this section we use reflection positivity and probabilistic arguments to provide basic estimates for some quantities of interest which are defined in the framework of the spin system. Most of our proofs however are formulated in the framework of the double dimer model.
In the whole section we use $G=(V,E)$ to denote the graph corresponding to the slab torus $\mathbb{S}_{L,K}$, we let $d$ be the degree of the origin and we omit the subscripts from $\langle\,\,\,\rangle_{L,K,\beta,h,r}$ when unnecessary for a lighter notation.
@@
Let $f_{1},f_{2}:[0,2\pi)\mapsto\mathbb{R}$ be two good functions which admit a Fourier series with a bounded numbers of terms. For any $\beta>0$, there exists $c=c(\beta,f_{1},f_{2})\in(0,\infty)$ such that, for any $h\in[0,1]$, $i,j\in\{1,2\}$, $L,K\in 2\mathbb{N}\cup\{1\}$, and disjoint sets $A,B\subset\mathbb{S}_{L,K}$
$\langle\prod_{x\in A}f_{1}(s_{x}^{i})\prod_{x\in B}f_{2}(s_{x}^{j})\rangle_{L,K,\beta,h,r}\leq c^{|A|+|B|}.$ (5.1)
@@
Using that by assumption for each $k\in\{1,2\}$ we can write $f_{k}(s_{x}^{i})=\sum_{n=-K}^{K}a_{n}e^{i\,n\,s_{x}^{i}}$ some $K<\infty$ and some real coefficients $a_{-K},\ldots a_{K}$ and calling $M=\max\{|a_{n}|\,:-K\leq n\leq K\}$ we obtain from Proposition 3.1 that
$\langle\prod_{x\in V}f_{k}(s^{i}_{x})\rangle_{L,K,\beta,h,r}=\sum\limits_{\boldsymbol{n}\in\{-K,\ldots,K\}^{V}}\langle\prod_{x\in V}a_{n_{x}}e^{i\,\,n_{x}^{i}\,s^{i}_{x}}\rangle\rangle_{L,K,\beta,h,r}$
$\leq{(2K+1)}^{|V|}\,\,{\max\{M,1\}}^{|V|}\max\Big{\{}\frac{1}{Z_{G,\beta,h,r}}\sum\limits_{\begin{subarray}{c}m\in\Sigma:\\ \partial m=(\boldsymbol{n}^{1},\,\,\boldsymbol{n}^{2})\end{subarray}}\tilde{\mu}_{\beta,h,r}(m)\,\,:\,\,\boldsymbol{n}=(\boldsymbol{n^{1}},\boldsymbol{n^{2}})\in\{-K,\ldots,K\}^{V\times V}\Big{\}}$
$\leq c^{|V|}{(2K+1)}^{|V|}\,\,{\max\{M,1\}}^{|V|}e^{2\beta d|V|+2\beta h|V|}.$
For the first inequality we used assumption that $U(n)\leq 1$ for each $n\in\mathbb{N}_{0}$. For the second inequality we summed over the number of dimers independently for each edge $e\in E\cup E_{g}$ (ignoring any constraint on their cardinality) and we used that $Z_{G,\beta,h,r}\geq c^{|V|}$, for some $c=c(\beta)\in(0,\infty)$, concluding the proof. ∎
@@
For any $\beta>0$, $\epsilon>0$, there exists a constant $c=c(\beta)>0$ and $M=M(\beta,\epsilon)<\infty$ such that for any $h\in[0,1]$, $r\in\{1,2\}$, $L,K\in 2\mathbb{N}\cup\{1\}$, $z\sim o$,
$\mathcal{P}_{L,K,\beta,h,r}\big{(}m_{\{o,z\}}^{i}>M\big{)}$ $\leq\epsilon,$ (5.3)
$\mathcal{P}_{L,K,\beta,h,r}\big{(}n_{o}^{i}>0\big{)}$ $\geq c.$ (5.4)

### v3

@@ 4 Reflection positivity and its consequences
In this section we show that the complex measure $\langle\,\,\,\rangle_{L,K,\beta,h,r}$ which has been introduced in Section 2 is reflection positive and use this property to derive convexity and monotonicity properties for the two-point function, following [33], and to derive a Chessboard estimate, which is a well-known consequence. The reflection positivity property of the complex measure reflects an equivalent reflection positivity property of the double-dimer model [50]. It relies on the complex duality relation between vertices having opposite parity in (2.5) and its proof is classical [55, Chapter 10].
@@ 4.1 Reflection positivity
We let $\mathcal{S}$ be a plane which is orthogonal to the Cartesian vector $e_{i}$ for some $i\in\{1,2,3\}$ and which intersects the midpoint of some edges.
Definition 4.1 (Reflection of sites, configurations and functions).
We let $\Theta:\mathbb{S}_{L,K}\mapsto\mathbb{S}_{L,K}$ be a reflection operator with respect to the plane $\mathcal{S}$.
By a slight abuse of notation we keep using $\Theta:\Omega_{s}\mapsto\Omega_{s}$ to denote a reflection operator that associates to each configuration $\boldsymbol{s}=(s_{x})_{x\in V}$ the reflected spin configuration $\Theta(\boldsymbol{s})=(s_{\Theta(x)})_{x\in\mathbb{S}_{L,K}}$.
@@
$\forall\boldsymbol{s}\in\Omega_{s}\qquad\Theta f(\boldsymbol{s}):={f(\Theta(\boldsymbol{s}))}.$
Definition 4.2 (Function domain).
We say that the function $f:\Omega_{s}\mapsto\mathbb{C}$ has domain $A\subset\mathbb{S}_{L,K}$ if for each pair of configurations $\boldsymbol{s},\boldsymbol{s}^{\prime}\in\Omega_{s}$ such that $s_{x}=s_{x}^{\prime}$ for each $x\in A$, $f(\boldsymbol{s})=f(\boldsymbol{s}^{\prime})$.
For example, the spin function $S_{x}^{k}$, $k=1,2$, defined in (2.1), has domain $\{x\}$.
Definition 4.3 (Torus halves and $\mathcal{A}^{\pm}$-functions ).
We denote by $\mathbb{S}_{L,K}^{\pm}$ two disjoint, connected, torus halves, $\mathbb{S}_{L,K}^{\pm}$, such that $\Theta(\mathbb{S}_{L,K}^{\pm})=\mathbb{S}_{L,K}^{\mp}$.
We let $\mathcal{A}^{\pm}$ be the set of functions $f:\Omega_{s}\mapsto\mathbb{C}$ which have domain $\mathbb{S}_{L,K}^{\pm}$.
Our spin variables satisfy the following identity, which is crucial for establishing reflection positivity of the spin measure: for each $x\in\mathbb{S}_{L,K}$,
[…]
The validity of (4.1) relies on the fact that the spin variables were defined in (2.1) with a complex chessboard property, see Remark 2.1. This condition further implies that the functions $\gamma_{z}$, appearing in the definition of the spin measure and being defined in (2.2), satisfy for each $x\in\mathbb{S}_{L,K}$,
$\Theta(\gamma_{x})=\overline{\gamma_{\Theta(x)}}.$ (4.2)
Given these key properties of the spin representation, the proof strategy is classical and can be found in several papers about reflection positivity, see for example [55, Chapter 10].
Proposition 4.4 (Reflection positivity).
@@
Let us define for any measurable function $f:\Omega_{s}\mapsto\mathbb{C}$ the operator
$\langle f\rangle_{0}:=\int_{\Omega_{s}}f(\boldsymbol{s})\,d\boldsymbol{s}.$
Being $\langle\,\,\rangle_{0}$ a product measure it is easy to see that, the measure is reflection positive, namely for $f,g\in\mathcal{A}^{+}$
$\langle f\overline{\Theta(f)}\rangle_{0}\geq 0,$ (4.4)
and
$\langle f\,\,\overline{\Theta(g)}\rangle_{0}=\langle\overline{g}\,\,{\Theta(f)}\rangle_{0},$ (4.5)
(see [55, Chapter 10]). We now define for any measurable function $f:\Omega_{s}\mapsto\mathbb{C}$ the operator
$\langle f\rangle_{1}:=\int_{\Omega_{s}}f(\boldsymbol{s})\,\boldsymbol{\gamma}(\boldsymbol{s})d\boldsymbol{s},$
where $\boldsymbol{\gamma}(\boldsymbol{s})=\prod_{z\in\mathbb{S}_{L,K}}\gamma_{z}(\boldsymbol{s})$ and $\gamma_{z}(\boldsymbol{s})$ has been defined in (2.2). Define also
$\boldsymbol{\gamma}^{\pm}(\boldsymbol{s})=\prod_{z\in\mathbb{S}^{\pm}_{L,K}}\gamma_{z}(\boldsymbol{s}).$
We see that $\boldsymbol{\gamma}^{\pm}\in\mathcal{A}^{\pm}$ and that, since (4.2) holds, $\Theta(\boldsymbol{\gamma}^{\pm})=\overline{\boldsymbol{\gamma}^{\mp}}.$ Hence, since $\langle\,\,\rangle_{0}$ is reflection positive, we deduce that also $\langle\,\,\rangle_{1}$ is reflection positive, namely for any $f,g\in\mathcal{A}^{+}$
$\langle f\overline{\Theta(f)}\rangle_{1}=\langle(f\boldsymbol{\gamma}^{+})\,\,\overline{\Theta\big(f\boldsymbol{\gamma}^{+}\big)}\rangle_{0}\geq 0,$ (4.6)
and
$\langle f\,\,\overline{\Theta(g)}\rangle_{1}=\langle(f\boldsymbol{\gamma}^{+})\,\,\overline{{\Theta\big(g\boldsymbol{\gamma}^{+}}}\big)\rangle_{0}=\langle{\overline{\big(g\boldsymbol{\gamma}^{+}}}\big)\,\,\Theta\big({f\boldsymbol{\gamma}^{+}}\big)\,\,\rangle_{0}=\langle\overline{g}\,\,{\Theta(f)}\rangle_{1}.$ (4.7)
We now prove that the operator $\langle\,\,\rangle_{L,K,\beta,h,r}$ is reflection positive. For this, let $E^{\pm}$ be the subset of edges of $\mathbb{S}_{L,K}$ which are entirely contained in $\mathbb{S}_{L,K}^{\pm}$ and $E_{R}$ be the set of edges intersecting both $\mathbb{S}_{L,K}^{-}$ and $\mathbb{S}_{L,K}^{+}$. We observe that we can write for any $f,g\in\mathcal{A}^{+}$
$\langle f\,\,\overline{\Theta g}\rangle_{L,K,\beta,h,r}=\langle f\,\,\overline{\Theta g}\,\,e^{H_{+}}e^{H_{-}}e^{\beta\sum_{\small\{x,y\}\in E_{R}}S_{x}^{1}\,\,\overline{\Theta(S_{x}^{1})}}e^{\beta\sum_{\small\{x,y\}\in E_{R}}S_{x}^{2}\,\,\overline{\Theta(S_{x}^{2})}}\rangle_{1}$ (4.8)
@@
$H_{\pm}(\boldsymbol{s}):=\beta\sum\limits_{\{i,j\}\in E^{\pm}}S_{i}^{1}S_{j}^{1}+\beta\sum\limits_{\{i,j\}\in E^{\pm}}S_{i}^{2}S_{j}^{2}+\,\beta\,h\sum\limits_{k\in\mathbb{S}_{L,K}^{\pm}}\,2\,\cos(rs^{1}_{k})$
and we used the convention that if $\{x,y\}\in E_{R}$, then $x\in\mathbb{S}_{L,K}^{+}$ and $y\in\mathbb{S}_{L,K}^{-}$. Noting that $e^{H_{\pm}}=\overline{\Theta e^{H_{\mp}}}$, expanding the last two exponentials in (4.8) we obtain
$\langle f\,\,\overline{\Theta g}\rangle_{L,K,\beta,h,r}\\ =\sum\limits_{m^{1},m^{2}\in{\mathbb{N}_{0}}^{E_{R}}}\Big\langle f\,\,e^{H_{+}}\,\overline{\Theta\Big(g\,e^{H_{+}}\,\Big)}\prod_{\{x,y\}\in E_{R}}\frac{\big(\beta\,S_{x}^{1}\,\,\,\overline{\Theta(S_{x}^{1}})\Big)^{m_{\{x,y\}}^{1}}}{m_{\{x,y\}}^{1}!}\frac{\big(\beta S_{x}^{2}\,\,\overline{\Theta(S_{x}^{2}})\Big)^{m_{\{x,y\}}^{2}}}{m^{2}_{\{x,y\}}!}\Big\rangle_{1}.$ (4.9)
We deduce from (4.6) that, for general measurable functions $f,g\in\mathcal{A}^{+}$, each term in the previous sum can be written as $\langle h\,\overline{\Theta(h^{\prime})}\rangle_{1}$ for some $h,h^{\prime}\in\mathcal{A}^{+}$. Using the symmetry property (4.7), we have $\langle h\,\overline{\Theta(h^{\prime})}\rangle_{1}=\langle\overline{h^{\prime}}\,\Theta(h)\rangle_{1}.$ Summing over all terms then yields the first claim of the proposition. Moreover, if $f=g$, we deduce that each term in the previous sum can be written as $\langle h\,\overline{\Theta(h)}\rangle_{1}$ for some $h\in\mathcal{A}^{+}$, and is then non-negative by (4.6). This gives the second claim and concludes the proof. ∎
@@ 4.2 Convexity and monotonicity of two-point functions
Reflection positivity is useful, as it allows one to derive monotonicity properties of two-point functions, such us the functions in the RHS of (3.16) and (3.17). Such functions can be expressed as averages of products of two components of a reflection invariant vector function, defined below.
Definition 4.5 (Reflection invariant vector function).
A family of functions
$F=(F_{x})_{x\in\mathbb{S}_{L,K}},\hskip 18.49988ptF_{x}:\Xi\to\mathbb{R},$
is called reflection invariant if for all $x,y,z\in\mathbb{S}_{L,K}$ the following hold:
@@
    $\Theta F_{x}=F_{\Theta x}$ for any reflection $\Theta$ through edges,
@@
    $\langle F_{x}F_{y}\rangle=\langle F_{x+z}F_{y+z}\rangle\geq 0$, where addition is understood with respect to the torus metric.
The next lemma will allow us to derive convexity properties for two-point functions.
@@
Consider any reflection plane $\mathcal{S}$, let $\Theta$ be the corresponding reflection operator and $\mathbb{S}_{L,K}^{\pm}\subset\mathbb{S}_{L,K}$ the corresponding torus halves as in Section 4.1. Let $(F_{x})_{x\in\mathbb{S}_{L,K}}$ be a reflection invariant vector of functions such that $F_{o}$ has domain $D\subset\mathbb{S}_{L,K}$, with $o\in D$, and define $\ell:=\sup\{d(o,x):x\in D\}$. Let $Q\subset\mathbb{S}_{L,K}$, $Q^{\pm}:=(Q\cap\mathbb{S}_{L,K}^{\pm})\cup\Theta(Q\cap\mathbb{S}_{L,K}^{\pm})$, assume that $|Q^{+}|=|Q^{-}|$ and that $d_{H}(Q^{\pm},\mathcal{S})>\ell$, where $d_{H}$ is the Hausdorff distance with respect to the $L^{2}$ norm. Then,
@@
$\mathcal{G}(x_{n})\geq\mathcal{G}(x_{n+2})\qquad\qquad\mathcal{G}(x_{k})\geq\mathcal{G}(x_{k+2}).$
For the proof of the proposition (given the analogous of Proposition 4.4) we refer to [33].
@@ 4.3 Chessboard estimates
[…]
@@
Let $f=(f_{t})_{t\in\mathbb{S}_{L,K}}$ be a sequence of real expectation functions having domain $\{o\}$. We have that,
@@
Given Proposition 4.4, the proof of Proposition 4.9 is classical and is based on performing a sequence of reflections, we refer for example to [55, Theorem 10.11] for its proof.
@@ 5 Probabilistic estimates
In this section we use reflection positivity and probabilistic arguments to provide basic estimates for some quantities of interest which are defined in the framework of the spin system. Most of our proofs however are formulated in the framework of the multi-occupancy double-dimer model.
In the whole section we use $G=(V,E)$ to denote the graph corresponding to the slab torus $\mathbb{S}_{L,K}$, we let $d$ be the degree of the origin (which might be $4$ or $6$), and we omit the subscripts from $\langle\,\,\,\rangle_{L,K,\beta,h,r}$ when unnecessary for a lighter notation. We assume that $L$ is even, thus ensuring that $\mathcal{D}_{G}\neq\emptyset$.
@@
Fix $\beta>0$, Let $f_{1},f_{2}:[0,2\pi)\mapsto\mathbb{R}$ be two functions which admit a Fourier series with a bounded numbers of terms and have all real-valued coefficients, namely there exists $R\in\mathbb{N}_{0}$ and some coefficients $a^{j}_{-R},\ldots a^{j}_{R}\in\mathbb{R}$ such that for each $s\in[0,2\pi)$ and $j\in\{1,2\}$
$f_{j}(s)=\sum_{n=-R}^{R}a^{j}_{n}e^{i\,n\,s}.$
Define
$M:=\max\big\{\max\{|a^{j}_{n}|\,:-R\leq n\leq R,j\in\{1,2\}\},\,1\big\}.$
Then, there exists $c=c(\beta)\in(0,\infty)$ such that, for any $h\in[0,1]$, $i,j\in\{1,2\}$, $L,K\in 2\mathbb{N}\cup\{1\}$, and disjoint sets $A,B\subset\mathbb{S}_{L,K}$
$\langle\prod_{x\in A}f_{1}(s_{x}^{i})\prod_{x\in B}f_{2}(s_{x}^{j})\rangle\leq(c\,M\,(R+1))^{|A|+|B|}.$ (5.1)
@@
Fix $i,j\in\{1,2\}$. We have
$\langle\prod_{x\in V}f_{j}(s^{i}_{x})\rangle=\sum\limits_{\boldsymbol{n}\in\{-R,\ldots,R\}^{V}}\langle\prod_{x\in V}a^{j}_{n_{x}}e^{i\,\,n_{x}\,s^{i}_{x}}\rangle$
$\leq{(2R+1)}^{|V|}\,\,{M}^{|V|}\max\Big\{\frac{1}{Z^{path}_{G,\beta,h,r}}\sum\limits_{\begin{subarray}{c}m\in\Sigma_{r}:\\ \partial m=(\boldsymbol{n}^{1},\,\,\boldsymbol{n}^{2})\end{subarray}}{\nu}_{\beta,h,r}(m)\,\,:\,\,\boldsymbol{n}=(\boldsymbol{n^{1}},\boldsymbol{n^{2}})\in\{-R,\ldots,R\}^{V\times V}\Big\}$
$\leq c^{|V|}{(2R+1)}^{|V|}\,\,{M}^{|V|}e^{2\beta|E|+2\beta h|V|}.$
For the first inequality we used Proposition 3.11 to express the average with respect to the spin measure as an average with respect to the measure of the multi-occupancy double dimer model (recall that, by Proposition 3.11, taking the average with respect to $(S_{x}^{i})^{n}$, which equals $e^{ins_{x}^{i}}$ or $e^{-ins_{x}^{i}}$, induces the constraint that there are $n$ dimers on $\{x,s\}$ of type $1$ or $2$). We then used the assumption that $U(n)\leq 1$ for each $n\in\mathbb{N}_{0}$, we observed that the number of possible matchings at $x$ is $n_{x}(w)!$, thus cancelling the factor $n_{x}(w)!$ in the denominator of (3.2), and we took the maximum over all possible ways to choose the number of dimers on each edge touching the source vertex. For the second inequality we used the definition of the weight $\nu_{\beta,h,r}$ (see Definition 3.6) and summed over the number of dimers independently for each edge $e\in E\cup E_{g}$ and each colour. We also used the fact that
$Z^{path}_{G,\beta,h,r}\geq c^{|V|},$ (5.3)
for some $c=c(\beta)\in(0,\infty)$. Indeed, since $G$ admits at least one dimer configuration, in order to lower bound $Z^{path}_{G,\beta,h,r}$ one can fix a dimer configuration and use the lower bound $Z^{path}_{G,\beta,h,r}\geq\mu_{\beta,h,r}(w)$, where $w\in\mathcal{W}^{\ell}$ is the configuration which has exactly $n$ loops of length two on each edge hosting a dimer, where $n$ is a fixed integer such that $U(n)>0$. This gives the desired lower bound on $Z^{path}_{G,\beta,h,r}$. ∎
@@
For any $\beta>0$, $\epsilon>0$, there exists a constant $c=c(\beta)>0$ and $M=M(\beta,\epsilon)<\infty$ such that for any $h\in[0,1]$, $r\in\{1,2\}$, $L,K\in 2\mathbb{N}\cup\{1\}$, $z,o\in V$ neighbours,
$P_{L,K,\beta,h,r}\big(m_{\{o,z\}}^{i}>M\big)$ $\leq\epsilon,$ (5.4)
$P_{L,K,\beta,h,r}\big(n_{o}>0\big)$ $\geq c.$ (5.5)